Propane, C3H8(g), and oxygen gas, O2(g), react to produce carbon dioxide, CO2(g), and water, H2O(g). If all gases are at 25.0oC and 102 kPa, what volume of carbon dioxide is produced when 5.44 g of propane reacts with excess oxygen?


a- 0.13 L

b- 1.05 L

c- 8.99 L

d- 65.6 L


1
Expert's answer
2021-04-21T04:04:27-0400

balanced equation is

C3H8+5O2→3CO2+4H2O

1 moles of propane produces 3 moles of CO2

Molar Mass of propane = 44.1

= 5.44/44.1 = 0.1234moles of propane

= 0.1234 × 3 = 0.3701moles of CO2

Molar Mass of CO2 = 44.01

= 44.01 × 0.3701

= 16.287g ofCO2